The December 31, 2015, balance sheet of Schism, Inc., showed long-term debt of $1,395,000, $139,000 in the common stock account and $2,640,000 in the additional paid-in surplus account. The December 31, 2016, balance sheet showed long-term debt of $1,570,000, $149,000 in the common stock account and $2,940,000 in the additional paid-in surplus account. The 2016 income statement showed an interest expense of $93,500 and the company paid out $144,000 in cash dividends during 2016. The firm’s net capital spending for 2016 was $950,000, and the firm reduced its net working capital investment by $124,000. What was the firm's cash flow to creditors during 2016? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars, e.g., 1,234,567.) Cash flow to creditors $ What was the firm’s cash flow to stockholders during 2016? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars, e.g., 1,234,567.) Cash flow to stockholders $ What was the firm’s cash flow from assets during 2016?
Answer #1

Cash Flow to creditors = Interest - New Debt Raised = 93,500 - (1570000- 1,395,000) = -81500
Cash flow to Shareholders = Dividends -(New Stocks issued) = 144,000 -( (2,940,000+149,000)-(2,640,000+139,000)) = -166,000

Cash flow from Assets = Cash Flow to Creditors + Cash Flow to Shareholders = -81500 -166,000 = -247,500

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well